OpenStudy (anonymous):

Solve the equation on the interval 0<(or equal to) theta<2pi a.) 2sin (theta)+4=5 b.)sin 3(theta)=1 c.) cos (2(theta)= -1/2 d.) 4sin(theta)+2=0

OpenStudy (misty1212):

HI!! one at a time dear

OpenStudy (anonymous):

\[0\le \theta<2\pi\]

OpenStudy (anonymous):

and oh sorry i just typed up the whole question

OpenStudy (misty1212):

it is 4 questions right?

OpenStudy (anonymous):

yeah but it was formatted like that so i just wrote the whole thing

OpenStudy (misty1212):

\[2\sin (\theta)+4=5\]

OpenStudy (misty1212):

first solve for \(\sin(\theta)\) by subtracting 4 and dividing by 2

OpenStudy (misty1212):

let me know when you get \[\sin(\theta)=\frac{1}{2}\]

OpenStudy (anonymous):

i got that

OpenStudy (misty1212):

then ask yourself if you know a number whose sine is one half

OpenStudy (misty1212):

or look on the unit circle

OpenStudy (anonymous):

ok isn't that pi/6?

OpenStudy (misty1212):

that is one of them yes

OpenStudy (misty1212):

then go across the unit circle and see that \(\frac{5\pi}{6}\) also works

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

others are similar

OpenStudy (anonymous):

hm ok ill try the next one on my own then and get back to you if i get stuck, thanks for your help! :)

OpenStudy (misty1212):

\[\sin(3\theta)=1\] ok that is a bit different

OpenStudy (anonymous):

oh ok sorry carry on

OpenStudy (misty1212):

first solve for \(3\theta\) then divide by 3

OpenStudy (misty1212):

\[\sin(3\theta)=1\\ 3\theta=\frac{\pi}{2}\\ \theta=\frac{\pi}{6}\] is one answer

OpenStudy (misty1212):

then go around the circle again \[\sin(3\theta)=1\\ 3\theta =\frac{5\pi}{2}\\ \theta=\frac{5\pi}{6}\]

OpenStudy (misty1212):

you got that? other two are like the first two

OpenStudy (anonymous):

where'd you get the 5pi/2 it's not even on the Unit Circle.. unless my eyes are just failing to see it aha

OpenStudy (anonymous):

oh wait it's not supposed to be on the unit circle, is it?

OpenStudy (misty1212):



OpenStudy (misty1212):

it is an angle the points on the unit circle are the coordinates

OpenStudy (misty1212):

the angle can go around and around and around that is why sine and cosine are periodic

OpenStudy (anonymous):

ohh ok